Calculate the present value, to the nearest cent, for each of the following:
a) Future value = $34000, Interest rate = 4% p.a, Time period = 4 years

Answer by jim_thompson5910(35256)


I'm assuming you're compounding the interest once per year.

The formula to use is
PV = FV/((1+r)^n)
where,
PV = present value = unknown
FV = future value = 34,000
r = 0.04 which is the decimal form of 4%
n = number of years = 4

So,
PV = FV/((1+r)^n)
PV = (34,000)/((1+0.04)^4)
PV = (34,000)/((1.04)^4)
PV = (34,000)/(1.16985856)
PV = 29,063.3424950107
PV = 29,063.34 dollars